OpenStudy (anonymous):

what is the sample size formula for ME=2√(P(1-P)/n) if ME= margin errors, P= percentage of people responding in a certain way, and n is the size of the sample?

OpenStudy (anonymous):

I need it to solve a problem......is it like this? n= (p(1-p)*2^2)/(ME^2)

OpenStudy (anonymous):

Is the given formula\[ME=2 \sqrt{\frac{ P(1-P) }{ n }}\]

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry, I was answering the other problems

OpenStudy (anonymous):

And is your answer\[n=\frac{ 4P(1-P) }{ ME^2 }\]

OpenStudy (anonymous):

well, I did it like this n= (p(1-p)*2^2)/(ME^2 )

OpenStudy (anonymous):

Right, but I'm having a hard time deciphering the brackets. Knowing that \(2^2=4\), is my last post the same as your answer?

OpenStudy (anonymous):

oh, jaja you are right. Didn't noticed that XD

OpenStudy (anonymous):

Well then, you are correct! Well done!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome